In 1980 the ratio of male students to female students at Frost College was 2 males to 3 females. Since then, the enrollment of male studnets in the college has increased by 400 and the enrollment of female students has remained the same. The ratio of ales to females is currently 1 to 1. How many students are currently enrolled at Frost College?
Answer by stanbon(48558)

In 1980 the ratio of male students to female students at Frost College was 2 males to 3 females.
Let # of male students be 2x; then # of female students was 3x
Total # of students in 1980 was 5x
--------------------------------
Since then, the enrollment of male studnets in the college has increased by 400 and the enrollment of female students has remained the same.
# of male is now 2x+400 ; # of female now is still 3x
Total number of students now is 5x+400
-------------------------
The ratio of ales to females is currently 1 to 1.
EQUATION:
(2x+400)/3x = 1/1
2x+400 = 3x
x = 400
------------------------------
How many students are currently enrolled at Frost College?
5x+400 = 5*400 + 500 = 2400 (this is current enrollment)
====================
Cheers,
Stan H.